OpenStudy (anonymous):

evaluate 4x-2y-z if x=2 y =3 and z -6

OpenStudy (anonymous):

8

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

so much for helping

OpenStudy (anonymous):

plug in the value of x, y and z u will get 4*2 -2*3 -(-6) = 8 - 6 +6 = 8

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

howu do this one x*x*x*x=?

OpenStudy (anonymous):

press shift then 8

OpenStudy (anonymous):

no that a question i have to answer